# Unraveling the Enigma of Double Descent: An In-depth Analysis through the Lens of Learned Feature Space

### v2

@@ Abstract
Double descent presents a counter-intuitive aspect within the machine learning domain, and researchers have observed its manifestation in various models and tasks. While some theoretical explanations have been proposed for this phenomenon in specific contexts, an accepted theory for its occurrence in deep learning remains yet to be established. In this study, we revisit the phenomenon of double descent and demonstrate that its occurrence is strongly influenced by the presence of noisy data. Through conducting a comprehensive analysis of the feature space of learned representations, we unveil that double descent arises in imperfect models trained with noisy data. We argue that double descent is a consequence of the model first learning the noisy data until interpolation and then adding implicit regularization via over-parameterization acquiring therefore capability to separate the information from the noise.
@@ 1 Introduction
The double descent phenomenon is prevalent across various machine learning models. In the context of linear models, a series of empirical and theoretical studies have emerged to analyze this phenomenon (Loog et al., 2020). In deep learning, a number of researchers have offered additional evidence supporting the presence and prevalence of the double descent phenomenon across a broader range of deep learning models, including CNNs, ResNets, and Transformers, in diverse experimental setups. Nakkiran et al. (2021) also reveals that this phenomenon is most pronounced when dealing with label noise in the training set, while can manifest even without label noise. Despite the widely accepted empirical observation of the double descent phenomenon within the domain of deep learning, comprehending the underlying mechanism of this occurrence in deep neural networks continues to pose a significant question.
@@
It is well-known that deep neural networks accomplish end-to-end learning through integrating feature learning with classifier training (Zhou, 2021). A neural network tailored for classification tasks can be divided into two components. The initial segment is dedicated to feature space transformation, specifically the conversion of the original feature space represented by the input layer to the final feature space represented by the final representation layer, in which a classifier is constructed. We refer to this final feature space as the learned feature space, which is shaped by the acquired representations within fully-trained neural networks. In this paper, we highlight the crucial role that noise plays in shaping the behavior of deep learning models through a comprehensive exploration and interpretation of the interplay among model width, noisy training data, and the learned feature space. We demonstrated that over-parameterized models when contrasted with under-parameterized ones, interpolate more correct training data around noisy ones of matching classes. This phenomenon demonstrates a correlation with the double descent phenomenon, and we propose that this is attributed to models acquiring the ability to effectively ‘isolate’ noise from the information in the training dataset. We posit that this phenomenon has not been previously documented and could serve to elucidate the internal mechanism of double descent.
@@
The main contributions of this study are summarized as follows. We proposed that there exists a robust correlation between the neural network’s interpolation strategy of noisy data and the double descent phenomenon. Utilizing the $k$-nearest neighbor algorithm, we indicate that this phenomenon results from broader architectures genuinely interpolating noisy samples among correct samples of the same original class, showcasing an enhanced capability to isolate noise information. We believe our investigation is able to shed light on the mechanism by which over-parameterization enhances generalization.
@@ 2 Related Work
The prevailing literature that explores the double descent phenomenon in deep neural networks learning provides a range of explanations: bias-variance decomposition (Yang et al., 2020); samples to parameters ratio (Belkin et al., 2020; Nakkiran et al., 2021); decision boundaries (Somepalli et al., 2022); and sharpless of interpolation (Gamba et al., 2022). While offering an extensive array of experimental findings that replicate the double descent phenomenon, Nakkiran et al. (2021) concluded that the emergence of double descent exhibits a robust correlation with the inclusion of noisy labels in the training dataset. Somepalli et al. (2022) further studied the decision boundaries of random data points and established a relationship between the fragmentation of class regions and the double descent phenomenon. Recently, Chaudhary et al. (2023) revisited the phenomenon and proposed that double descent may not inherently characterize deep learning.
@@
In contrast to existing studies, our approach distinguishes itself by exploring the phenomenon through the lens of the learned feature space generated by trained models. The most related works to ours are the concurrent ones by Gamba et al. (2022), which measure sharpness at clean and noisily-labeled data points and empirically showed that smooth interpolation emerging both for large over-parameterized networks and trained large models confidently predicts the noisy training targets over large volumes around each training point. While their study also entailed an analysis involving separate evaluations of clean and noisily labeled data points, we employed the $k$-nearest neighbor algorithm to infer the inter-relationship between these two categories of training data in the learned feature space and demonstrated that for over-parameterized models, noisy labeled data tends to be classified alongside its genuine counterparts. Our approach suggests that the double descent phenomenon is a consequence of deep learning models learning how to separate the noise from the information in the training set.
@@ 3.1 Experiment Setup
  • Fully Connected Neural Networks (SimpleFC): We adopt a simple two-layer fully connected neural network with varying width $k$ on the first hidden layer, for $k$ in the range of [1, 1000]. The second layer is adopted as the classifier.
  • Standard Convolutional Networks (CNNs): We consider a family of standard convolutional neural networks formed by 4 convolutional stages of controlled base width [$k$, $2k$, $4k$, $8k$], for $k$ in the range of [1, 64], along with a fully connected layer as the classifier. The MaxPool is [2, 2, 2, 4]. For all the convolution layers, the kernel size = 3, stride = 1, and padding = 1. This architecture implementation is adopted from Nakkiran et al. (2021).
  • Residual Neural Networks (ResNet18s): We parameterize a family of ResNet18s from He et al. (2016) using 4 ResNet blocks, each consisting of two BatchNorm-ReLU-convolution layers. We scale the layer width (number of filters) of convolutional layers as [$k$, $2k$, $4k$, $8k$] for a varied range of $k$ within [1, 64] and the strides are [1, 2, 2, 2]. The implementation is adopted from https://github.com/kuangliu/pytorch-cifar.git.
All networks are trained using the Stochastic Gradient Descent (SGD) optimizer with Cross-Entropy Loss, zero momentum, and decreasing learning rate. A batch size of 128 is employed, and no explicit data augmentation or dropout techniques are utilized in our experimental setup. Utilizing a sequence of models trained under a uniform training scheme showcases the double descent phenomenon across different models. We refer our reader to Appendix A for further information regarding our experiment setup and scaling of model parameterization with layer width k.
@@ 3.2 Introduction of Label Noise
While the presence of label errors has been established in the majority of commonly used benchmarking datasets including MNIST and CIFAR-10 (Zhang, 2017; Northcutt et al., 2021b), for the purpose of facilitating comparative analysis of our findings, we intentionally introduce explicit label noise. In our experimental setup, label noise with probability $p$ signifies that during training, samples are assigned a uniformly random label with probability $p$, while possessing the correct label with probability $(1-p)$ otherwise. We interpret a neural network with input variable $x\in R^{d}$ and learnable parameter $\theta$, incorporating all weights and biases. The training dataset is characterized as a collection of training points $(x_{n},y_{n})$, where $n$ ranges from 1 to $N$. In a training dataset with $N$ data points, it comprises $m=p\times N$ noise labeled data points and $\hat{m}=(1-p)\times N$ clean data points.
@@
It is important to note that the label noise is sampled only once and not per epoch, and this original and random label information is stored before each experiment. Each experiment is replicated multiple times, and the final result is obtained by averaging the outcomes. Owing to the presence of inherent label errors, the probability $p$ is solely employed to signify the extent of explicit label noise and does not reflect the overall noise level present in the dataset.
3.3 Learned Feature Space Inference Strategy
@@
Our primary research inquiry aims to gain insights into the diverse strategies employed by deep neural networks of various widths in interpolating noisy labeled data in classification tasks. Based on the existence of benign over-parameterization, we assume that test images, akin to training images mislabeled, are more likely to be correctly classified by over-parameterized models. For instance, we anticipate that an optimal classifier should yield accurate predictions for unseen images, even if it was trained on similar images with an adversarial label. Thus, we further hypothesize that the closest neighbors of mislabeled training images are correctly classified. To validate our hypothesis, we adopted the $k$-nearest neighbour algorithm with cosine similarity to infer the relative locations of clean and noisy labeled data in the learned feature space. This methodology allows us to delineate the prediction strategy through an indirect approach and subsequently compare our findings with the overall generalization performance of these pre-trained models.
@@
We consider L-layer neural networks, while representations on each layer are represented by $f_{l}(X)$, for $l=[1,...,L]$. Our investigation centered on the penultimate layer representations before classification denoted as $f_{L-1}(X)$. Acknowledging the difficulties of inferring high-dimensional feature spaces, our approach mainly entails characterizing the relative positioning of training data points. This involves establishing a connection between the hidden features of noisy data points in the learned feature space and their $k$-nearest neighbors within the complementary subset of clean data. So we calculate the prediction accuracy $P$ of mislabeled training data and the majority of its nearest neighbours in feature space are in the same class:
@@
In this context, $(x_{1},y_{1}),\ldots,(x_{m},y_{m})$ denotes the subset of noisy training data with original labels, and $(x_{i,1},y_{i,1}),\ldots,(x_{i,\hat{m}},y_{i,\hat{m}}))$ denotes the subset of clean training data as neighbours to $x_{i}$ for all $i=[1,\ldots,m]$. Given the cosine similarity $S_{C}(A,B):=\frac{A\cdot B}{\|A\|\|B\|}$ between two representations A and B, let $f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,\hat{m}})$ be a reordering of the representations of the clean training data such that $S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,1}))\leq\cdots\leq S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,\hat{m}}))$. The closest $k$ neighbours are selected, and a Majority function $M(f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,k}))$ describes the procedure for making predictions based on the majority label among the k-nearest neighbors. The predictions are subsequently compared with the original labels of noisy data, and $P$ denotes the proportion of accurate predictions among all the noisy data.
In line with our hypothesis, the prediction accuracy $P$ is indicative of the neural architecture’s interpolation approach toward these noisy labeled data points. Due to the uncertain distribution of data and their representations, the $k$-NN method functions solely as an estimation technique, and our comparison focuses on variations in $P$.
In the next section, we present our empirical study and results of the learned feature space inference for trained neural networks of various widths in relationship to double descent.
@@ 4.1 MNIST
‘ (a) SimpleFC on MNIST $(p=0\%)$ (b) SimpleFC on MNIST $(p=10\%)$ (c) SimpleFC on MNIST $(p=20\%)$
@@
Figure 1: The phenomenon of double descent on two-layer fully connected neural networks trained on MNIST $(N=4000)$, under varying explicit label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labeled data denoted as $P$ when $p>0$. Optimized with SGD for 4000 epochs and decreasing learning rate. The test error curve of $p=[10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
@@
We commence our exploration by revisiting one of the earliest experiments that revealed the presence of the double descent phenomenon, as presented in the work of Belkin et al. (2019). This experiment involved the training of a basic two-layer fully connected neural network on a dataset comprising 4000 samples from the MNIST dataset ($N=4000$). When zero label noise ($p=0\%$) is introduced, as depicted in Figure.1(a), both the train and the test error curve decrease monotonically with increasing model width. The training error exhibits a faster decline and approaches zero at approximately $k=15$, indicating the occurrence of an interpolation threshold under these conditions. The test error still decreases after this interpolation threshold, demonstrating a benign over-parameterization. This noticeable alteration in the shape of the test error curve at the interpolation threshold may be attributed to the intrinsic label errors of MNIST (Northcutt et al., 2021a).
@@
With label noise ($p=10\%/20\%$) introduced in the training dataset, a more distinguishable test error peak starts to form around the interpolation threshold as shown in Figure.1(b, c). Starting with small models, the test error curve exhibits a U-shaped pattern in the first half and the bias-variance trade-off is reached at the bottom of the U-shaped curve. The decrease of both train and test error before this bias-variance trade-off indicates the model undergoes an under-fitting stage. With the expansion of the model size, the model transitions into an overfitting regime, leading to a peak in the test error around the interpolation threshold, while the train error converges to 0. Finally, after surpassing this test error peak, the testing loss experiences a decline once more, showcasing how highly over-parameterized models can ultimately yield improved generalization performance. This intriguing behavior is referred to as double descent. As the test loss undergoes a pattern of first decreasing, then increasing, and finally decreasing again, the test accuracy exhibits the opposite trend, i.e., first increasing, then decreasing, and then increasing once more. As Nakkiran et al. (2021) reported, the test loss peak is positively correlated with the label noise ratio $p$. In other words, as the label noise ratio increases, the peak in the test loss also tends to increase accordingly.
@@
To comprehend the influence of noisy labeled data on the variation of the learned model during the double descent phenomenon, we examine the percentage $P$ of noisy data, and the majority of its nearest neighbours in feature space are in the same class. In this experiment setup, as depicted in Figure.1(b, c), the trajectory of the $P$ curve is in accordance with the pattern observed in the test accuracy curve, indicating a consistent alignment between the two. When test loss first decreases as the model expands, test accuracy and $P$ both rise. With the introduction of label noise, test accuracy decreases while $P$ fluctuates around 70%. When passing the interpolation threshold, both the test accuracy and $P$ significantly rise to 85% and nearly 100%, substantially affirming the validity of our hypothesis. In the meantime, the trend of test loss is opposite to both test accuracy and $P$, while changing correspondingly before and after the interpolation threshold.
@@
Recall that the noisy data points will be predicted based on their randomly assigned labels during this stage as the model fully interpolates the training set, we interpret the ascending $P$ curve in the over-parameterized region as a valid ‘isolation’ of the noisy labeled data in the learned feature space. When the neural network positions the learned representation of noisy data points closer to the class associated with their ‘incorrect’ labels, it results in lower $P$ prediction accuracy. Under these circumstances, similar test samples also tend to be predicted incorrectly due to the influence of these noisy labels. If the neural network successfully identifies the underlying data patterns and interpolates the correct samples around noisy data points, the prediction accuracy $P$ will be high by nearby class members. Consequently, test samples resembling the noisy training data have a higher likelihood of being accurately identified due to their proximity to the interpolated surrounding clean data points and becoming less susceptible to the influence of the ‘isolated’ noisy labeled data. This explanation accounts for the improved generalization performance for over-parameterized models.
4.2 CIFAR-10
@@
Figure 2: The phenomenon of double descent on five-layer CNNs trained on CIFAR-10 ($N=50000$), under varying label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labeled data denoted as $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve of $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
After a baseline case of deep learning on MNIST, we extended the experiment result to image classification on CIFAR-10, using two famous neural network architectures: CNN and ResNet. We first visit the experiment when the complete CIFAR-10 dataset ($N=50000$) with zero explicit label noise ($p=0\%$) is trained with five-layer CNNs: we observe that in Figure.2(a), a proposed double descent peak arises around the interpolation threshold. We posit that the variance in observations could stem from a higher ratio of label errors in CIFAR-10 compared to MNIST, as suggested by research estimates (Zhang, 2017; Northcutt et al., 2021b). Nevertheless, when additional label noise ($p=10\%/20\%$) is introduced, the test error peak still increases accordingly, proving the positive correlation between noise and double descent according to Figure.2(b,c). While label noise increases the difficulties of models fitting the train set and shifts the interpolation threshold rightwards, the test error peak also shifts correspondingly.
When we review the noisy labeled data prediction test in the learned feature space, we can still observe a correlation between the trend of generalization performance and prediction accuracy $P$. When test accuracy first increases and then decreases as the model expands, $P$ rises to about 60% and then decreases before the threshold when the peak of test error arises. Then, as the test error reduces and test accuracy rises, the $P$ curve rises again to about 60% for over-parameterized models. The observed correlation between the trend of test loss and $P$ also remains consistent. The same conclusion can be derived from the experiment context Figure.3 of ResNet18s trained on the sample task.
In conclusion, by constructing $P$ to represent the distribution of noisy labeled training data in the learned feature space, we propose that this k-NN prediction ability of the model is connected to the size of the model in the same way with double descent. We suggest that this correlation can be attributed to the model’s interpolation of correct samples around the noise labels, effectively isolating the misleading effect of the latter.

### v3

@@ Abstract
Double descent presents a counter-intuitive aspect within the machine learning domain, and researchers have observed its manifestation in various models and tasks. While some theoretical explanations have been proposed for this phenomenon in specific contexts, an accepted theory for its occurring mechanism in deep learning remains yet to be established. In this study, we revisit the phenomenon of double descent and demonstrate that the presence of noisy data strongly influences its occurrence. By comprehensively analysing the feature space of learned representations, we unveil that double descent arises in imperfect models trained with noisy data. We argue that while small and intermediate models before the interpolation threshold follow the traditional bias-variance trade-off, over-parameterized models interpolate noisy samples among robust data thus acquiring the capability to separate the information from the noise. The source code is available at https://github.com/Yufei-Gu-451/double_descent_inference.git.
@@ 1 Introduction
The double descent phenomenon is prevalent across various machine learning models. In the context of linear models, a series of empirical and theoretical studies have emerged to analyze this phenomenon (Bartlett et al., 2020; Loog et al., 2020; Belkin et al., 2020; Deng et al., 2022; Hastie et al., 2022). In deep learning, many researchers have offered additional evidence supporting the presence and prevalence of the double descent phenomenon across a broader range of deep learning models, including CNNs, ResNets, and Transformers, in diverse experimental setups (Nakkiran et al., 2021; Chang et al., 2021; Liu et al., 2022; Gamba et al., 2022). Nakkiran et al. (2021) specifically concluded that this phenomenon is most pronounced when dealing with label noise in the training set, while can manifest even without label noise for the first time. Despite the widely accepted empirical observation of the double descent phenomenon within the domain of deep learning, comprehending the underlying mechanism of this occurrence in deep neural networks continues to pose a significant question.
It is well-known that deep neural networks accomplish end-to-end learning frameworks through integrating feature learning with classifier training (Zhou, 2021). A neural network tailored for classification tasks can be divided into two components. The initial segment is dedicated to a feature space transformation process by selecting information pertinent to the output while discarding information associated with the input. Specifically, the conversion of the original feature space represented by the input layer to the final feature space represented by the final representation layer is referred to as the learned feature space, which is shaped by the acquired representations within fully-trained neural networks. The subsequent stage typically entails constructing a classifier based on this learned feature space, intending to complete machine-learning tasks.
@@
In this paper, we highlight the crucial role that noisy samples play in shaping the behaviour of deep learning models through a comprehensive exploration of the interplay among model width, noisy training data, and interpretation of the learned feature space. We demonstrated that while small and intermediate models follow the traditional bias-variance trade-off, over-parameterized models which passed the interpolation threshold, interpolate more correct training data around noisy ones of matching classes. This phenomenon demonstrates a correlation with the double descent phenomenon, and we propose that this is attributed to models acquiring the ability to effectively ‘isolate’ noise from the information in the training dataset. We posit that this phenomenon has not been previously documented and could serve to elucidate the internal mechanism of double descent.
@@
The main contributions of this study are summarized as follows. We proposed that there exists a robust correlation between the neural network’s interpolation strategy of noisy data and the double descent phenomenon. Utilizing the $k$-nearest neighbour algorithm, we indicate that this phenomenon results from broader architectures genuinely interpolating noisy samples among correct samples of the same original class, showcasing an enhanced capability to isolate noise information. Through conducting experiments on various neural architectures trained on MNIST and CIFAR-10, we demonstrated that the label prediction accuracy of noisy samples strongly aligns with the test performance on unseen data. We believe our investigation can shed light on the mechanism by which over-parameterization enhances generalization.
@@ 2 Related Work
The prevailing literature that explores the double descent phenomenon in deep neural networks learning provides a range of explanations: bias-variance decomposition (Yang et al., 2020); samples to parameters ratio (Belkin et al., 2020; Nakkiran et al., 2021); decision boundaries (Somepalli et al., 2022); and the sharpless of interpolation (Gamba et al., 2022). While offering an extensive array of experimental findings that replicate the double descent phenomenon, Nakkiran et al. (2021) concluded that the emergence of double descent exhibits a robust correlation with the inclusion of noisy labels in the training dataset. Somepalli et al. (2022) further studied the decision boundaries of random data points and established a relationship between the fragmentation of class regions and the double descent phenomenon. Recently, Chaudhary et al. (2023) revisited the phenomenon and proposed that double descent may not inherently characterize deep learning.
@@
In contrast to existing studies, our approach distinguishes itself by exploring the phenomenon through the lens of the learned feature space generated by trained models. The most related works to ours are the concurrent ones by Gamba et al. (2022), which measure sharpness at clean and noisily-labeled data points and empirically showed that smooth interpolation emerging both for large over-parameterized networks and trained large models confidently predicts the noisy training targets over large volumes around each training point. While their study also entailed an analysis involving separate evaluations of clean and noisily labelled data points, we employed the $k$-nearest neighbour algorithm to infer the inter-relationship between these two categories of training data in the learned feature space and demonstrated that for over-parameterized models, noisy labelled data tends to be classified alongside its genuine counterparts. Our approach suggests that the double descent phenomenon is a consequence of deep learning models learning how to separate the noise from the information in the training set.
@@ 3.1 Experiment Setup
  • Fully Connected Neural Networks (FCNN): We adopt a simple two-layer fully connected neural network with varying width $k$ on the first hidden layer, for $k$ in the range of [1, 1000]. The second fully connected layer is adopted as the classifier.
  • Standard Convolutional Networks (CNN): We consider a family of standard convolutional neural networks formed by 4 convolutional stages of controlled base width [$k$, $2k$, $4k$, $8k$], for $k$ in the range of [1, 64], along with a fully connected layer as the classifier. The MaxPool is [2, 2, 2, 4]. For all the convolution layers, the kernel size = 3, stride = 1, and padding = 1. This architecture implementation is adopted from Nakkiran et al. (2021).
  • Residual Neural Networks (ResNet18): We parameterize a family of ResNet18s from He et al. (2016) using 4 ResNet blocks, each consisting of two BatchNorm-ReLU-convolution layers. We scale the layer width (number of filters) of convolutional layers as [$k$, $2k$, $4k$, $8k$] for a varied range of $k$ within [1, 64] and the strides are [1, 2, 2, 2]. The implementation is adopted from https://github.com/kuangliu/pytorch-cifar.git.
We describe the experiment setup and hyper-parameters we used in training all neural architectures below. All networks are trained using the Stochastic Gradient Descent (SGD) optimizer with zero momentum and Cross-entropy Loss for evaluation. A separate learning rate scheme is employed for different models: $lr=\frac{0.05}{\sqrt{1+[epoch/50]}}$ for FCNNs and $lr=\frac{0.05}{\sqrt{1+[epoch*10]}}$ for CNNs and ResNet18s. For every model, the learning rate starts with an initial value of 0.05 and updates every 50 epochs. A batch size of 128 is employed, and no explicit data augmentation or regularization techniques (including dropouts or pruning) are utilized in our experimental setup. Utilizing a sequence of models trained under a uniform training scheme showcases the double descent phenomenon across different models.
@@ 3.2 Introduction of Label Noise
While the presence of label errors has been established in the majority of commonly used benchmarking datasets including MNIST and CIFAR-10 (Zhang, 2017; Northcutt et al., 2021b), to facilitate comparative analysis of our findings, we intentionally introduce explicit label noise. In our experimental setup, label noise with probability $p$ signifies that during training, samples are assigned a uniformly random label with probability $p$, while possessing the correct label with probability $(1-p)$ otherwise. We interpret a neural network with input variable $x\in R^{d}$ and learnable parameter $\theta$, incorporating all weights and biases. The training dataset is characterized as a collection of training points $(x_{n},y_{n})$, where $n$ ranges from 1 to $N$. A training dataset with $N$ data points comprises $m=p\times N$ noise labelled data points and $\hat{m}=(1-p)\times N$ clean data points.
@@
It is important to note that the label noise is sampled only once and not per epoch, and this original and random label information is stored before each experiment. Each experiment is replicated multiple times, and the final result is obtained by averaging the outcomes. Owing to the presence of inherent label errors, the probability $p$ is solely employed to signify the extent of explicit label noise and does not reflect the overall noise level present in the dataset. Our primary experimental conclusion is drawn from comparing results obtained under a dataset with a relatively noisy level.
3.3 Learned Feature Space Interpretation Strategy
@@
Our primary research inquiry aims to gain insights into the diverse strategies employed by deep neural networks of various widths in interpolating noisy labelled data in classification tasks. Based on the existence of benign over-parameterization, we assume that test images, akin to training images mislabeled, are more likely to be correctly classified by over-parameterized models. For instance, we anticipate that an optimal classifier should yield accurate predictions for unseen images, even if it were trained on similar images with an adversarial label. Thus, we further hypothesize that the closest neighbours of mislabeled training images are correctly classified for a model with better generalization performance. To validate our hypothesis, we adopted the $k$-nearest neighbour algorithm with cosine similarity to interpret the relative locations of clean and noisy labelled data in the learned feature space. This methodology allows us to delineate the prediction strategy indirectly and subsequently compare our findings with the overall generalization performance of these pre-trained models.
@@
We consider L-layer neural networks, while representations on each layer are represented by $f_{l}(X)$, for $l=[1,...,L]$. Our investigation centred on the penultimate layer representations before classification denoted as $f_{L-1}(X)$. Acknowledging the difficulties of interpreting high-dimensional feature spaces, our approach mainly entails characterizing the relative positioning of training data points. This involves establishing a connection between the hidden features of noisy data points in the learned feature space and their $k$-nearest neighbours within the complementary subset of clean data. So we calculate the prediction accuracy $P$ of mislabeled training data and the majority of its nearest neighbours in feature space are in the same class:
@@
In this context, $(x_{1},y_{1}),\ldots,(x_{m},y_{m})$ denotes the subset of noisy training data with original labels, and $(x_{i,1},y_{i,1}),\ldots,(x_{i,\hat{m}},y_{i,\hat{m}}))$ denotes the subset of clean training data as neighbours to $x_{i}$ for all $i=[1,\ldots,m]$. Given the cosine similarity $S_{C}(A,B):=\frac{A\cdot B}{\|A\|\|B\|}$ between two representations A and B, let $f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,\hat{m}})$ be a reordering of the representations of the clean training data such that $S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,1}))\leq\cdots\leq S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,\hat{m}}))$. The closest $k$ neighbours are selected, and a Majority function $M(f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,k}))$ describes the procedure for making predictions based on the majority label among the k-nearest neighbours. The predictions are subsequently compared with the original labels of noisy data, and $P$ denotes the proportion of accurate predictions among all the noisy data.
In line with our hypothesis, the prediction accuracy $P$ is indicative of the neural architecture’s interpolation approach toward these noisy labelled data points. Due to the uncertain distribution of data and their representations, the $k$-NN method functions solely as an estimation technique, and our comparison focuses on variations in $P$. In the next section, we present our empirical study and results of the learned feature space interpretation for trained neural networks of various widths in relationship to double descent.
@@ 4 Experiments
‘ (a) FCNN on MNIST $(p=0\%)$ (b) FCNN on MNIST $(p=10\%)$ (c) FCNN on MNIST $(p=20\%)$
@@
Figure 1: The phenomenon of double descent on two-layer FCNNs trained on MNIST $(N=4000)$, under varying explicit label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labelled data denoted as $P$ when $p>0$. Optimized with SGD for 4000 epochs and decreasing learning rate. The test error curve of $p=[10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
@@
We commence our exploration by revisiting one of the earliest experiments that revealed the presence of the double descent phenomenon, as presented in the work of Belkin et al. (2019). This experiment involved the training of a basic two-layer FCNN on a dataset comprising 4000 samples from the MNIST dataset ($N=4000$). When zero label noise ($p=0\%$) is introduced, as depicted in Figure.1(a), both the train and the test error curve decrease monotonically with increasing model width. The training error exhibits a faster decline and approaches zero at approximately $k=15$, indicating the occurrence of an interpolation threshold under these conditions. The test error continues to decrease after this interpolation threshold, demonstrating a benign over-parameterization. This noticeable alteration in the shape of the test error curve at the interpolation threshold may be attributed to the intrinsic label errors of MNIST (Northcutt et al., 2021a).
@@
With external label noise ($p=10\%/20\%$) introduced in the training dataset, a more distinguishable peak starts to form in the test error around the interpolation threshold as shown in Figure.1(b, c). Starting with small models, the test error curve exhibits a U-shaped pattern in the first half and the bias-variance trade-off is reached at the bottom. The decrease of both train and test error before this bias-variance trade-off indicates the model undergoes an under-fitting stage. With the expansion of the model size, the model transitions into an overfitting regime, leading to a peak in the test error around the interpolation threshold, while the train error converges to 0. Finally, after surpassing this test error peak, the testing loss experiences a decline once more, showcasing how highly over-parameterized models lead to improved generalization performance. This intriguing behaviour is referred to as double descent. Meanwhile, the test accuracy exhibits the opposite trend to the test loss. As Nakkiran et al. (2021) reported, the test loss peak is positively correlated with the label noise ratio $p$. In other words, as the label noise ratio increases, the peak in the test loss also tends to increase accordingly.
@@
To comprehend the influence of noisy labelled data on the variation of the learned model during the double descent phenomenon, we examine the percentage $P$ of noisy data, and if the majority of its nearest neighbours in feature space are in the same class. As depicted in Figure.1(b, c), the trajectory of the $P$ curve follows the pattern observed in the test accuracy curve, indicating a consistent alignment between the two. When test loss first decreases as the model expands, test accuracy and $P$ both rise. With the introduction of label noise, test accuracy decreases while $P$ fluctuates around 70%. When passing the interpolation threshold, both the test accuracy and $P$ significantly rise to 85% and nearly 100%, substantially affirming the validity of our hypothesis. In the meantime, the trend of test loss is opposite to both test accuracy and $P$, while changing correspondingly before and after the interpolation threshold. The computation of percentage $P$ relies solely on the training dataset and does not incorporate unseen data. Consequently, this percentage $P$ may be regarded as a weak predictor of generalization performance, though it necessitates knowledge of the distribution between clean and noisy data.
@@
Recall that the noisy data points will be predicted based on their randomly assigned labels during this stage as the model fully interpolates the training set, we interpret the ascending $P$ curve in the over-parameterized region as a valid ‘isolation’ of the noisy labelled data in the learned feature space. Consequently, test samples resembling the noisy training data have a higher likelihood of being accurately identified due to their proximity to the interpolated surrounding clean data points and becoming less susceptible to the influence of the ‘isolated’ noisy labelled data. This explanation accounts for the improved generalization performance for over-parameterized models.
@@
Figure 2: The phenomenon of double descent on five-layer CNNs trained on CIFAR-10 ($N=50000$), under varying label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labelled data denoted as $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve of $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
After a baseline case of learning with FCNNs on MNIST, we extended the experiment results to image classification on the CIFAR-10 dataset, using two famous neural network architectures: five-layer CNN and ResNet18. We first visit the experiment when CIFAR-10 ($N=50000$) with zero explicit label noise ($p=0\%$) is trained with CNNs: Figure 2(a) demonstrated a proposed double descent peak arises around the interpolation threshold. We posit that the variance in observations could stem from a higher ratio of label errors in CIFAR-10 compared to MNIST, as suggested by research estimates (Zhang, 2017; Northcutt et al., 2021b). When additional label noise ($p=10\%/20\%$) is introduced in Figure 2(b,c), the test error peak increases accordingly, proving the positive correlation between noise and double descent. While label noise increases the difficulties of models fitting the train set and shifts the interpolation threshold rightwards, the test error peak also shifts correspondingly.
When we review the noisy labelled data prediction test in the learned feature space, we can also observe an aligned correlation between the trend of generalization performance and prediction accuracy $P$. When test accuracy first increases and then decreases as the model expands, $P$ rises to about 60% and then decreases before the threshold when the peak of test error arises. Then, as the test error reduces and test accuracy rises, the $P$ curve rises again to about 60% for over-parameterized models. The observed correlation between the trend of test loss and $P$ remains consistent. An important observation is that the prediction accuracy $P$ demonstrated in the experiments of CNNs is lower compared to those on MNIST. We attribute this variance to the intricate feature space representations, which are more challenging to capture using cosine similarities. Our analysis focused more on how the interpolation strategy varied across the same neural architecture of different widths, and the proposed k-NN methodology only served as an interpretation technique.